OpenStudy (anonymous):

If you are finding the line perpendicular to x + 3y = 5, what is the slope of the perpendicular line? A. -3/5 B. 1/3 C. 3 D. -3

OpenStudy (anonymous):

x + 3y = 5 3y = 5 - x y = 5/3 - x/3 this is the slope intercept form , where x's coefficeint is the slope

OpenStudy (anonymous):

so -1/3 is the slope of this line for a line perpendicular to this , the slope is -1/slope of line

OpenStudy (anonymous):

= -1/-1/3 = 3

OpenStudy (jdoe0001):

the slope of perpendicular lines are NEGATIVE RECIPROCAL so if you have a line with slope "a" for one, the perpendicular line to that will have a slope of RECIPROCAL of that 1/a NEGATIVE of that - 1/a

OpenStudy (mathlegend):

Do you understand @musa19965 ?

OpenStudy (anonymous):

Kinda